In July, the one year interest rate is 12.63% on British pounds and 9.13% on U.S. dollars. The current exchange rate is USD1.63/GBP. Suppose a change in expectations regarding future U.S. inflation causes the expected future spot rate to decline to $1.49/GBP. Let's assume that the British interest rate stayed at 12.63% because there has been no change in expectations of British inflation.

What would the U.S. interest rate expected to be?
Answer #1

Future spot rate = Current spot rate*(1+Ru)/(1+Rb)

Ru = Interest rate in US

Rb = Interest rate In British

1.49 = 1.63*(1+Ru)/(1+0.1263)

(1.49*1.1263) = 1.63*(1+Ru)

1.6782 = 1.63*(1+Ru)

1.0296 = 1+Ru

Ru = 0.0296

That means expected Interest rate in USA will be = 2.96%.

I hope my efforts will be fruitful to you....?

​​​​​​